# Dynamics in an emergent quantum-like state space generated by a nonlinear classical network

### v3

@@ Abstract
This work exploits a framework whereby a graph (in the mathematical sense) serves to connect a classical system to a state space that we call ‘quantum-like’ (QL). The QL states comprise arbitrary superpositions of states in a tensor product basis. The graph plays a special dual role by directing design of the classical system and defining the state space. We study a specific example of a large, dynamical classical system—a system of coupled phase oscillators—that maps, via a graph, to the QL state space. We investigate how mixedness of the state diminishes or increases as the underlying classical system synchronizes or de-synchronizes respectively. This shows the interplay between the nonlinear dynamics of the variables of the classical system and the QL state space. We prove that maps from one time point to another in the state space are linear maps. In the limit of a strongly phase-locked classical network—that is, where couplings between phase oscillators are very large—the state space evolves according to unitary dynamics, whereas in the cases of weaker synchronization, the classical variables act as a hidden environment that promotes decoherence of superpositions. We examine how similar the properties of QL states are to quantum states. We find that during decoherence of the QL states, the off-diagonal density matrix elements decay and that this decay can be observed in any basis we choose for measurement. More surprisingly, we show that a no-cloning theorem (that is, a state of a QL bit cannot be copied) applies not only to the QL states, but also to the underlying classical system.
@@ I Introduction
In recent work we have described how a classical system can be mapped, via a graph representation, to a state space that has many properties similar to a quantum state space[1, 2]. For that reason, we call it a quantum-like (QL) state space, and refer to the states as QL states. The graphs needed as a basis for these constructions are called QL bit graphs (or simply QL bits). By applying a product operation to QL bit graphs we can produce graphs whose states comprise arbitrary (controllable) superpositions of states in a tensor product basis—entangled states. These same graphs can also be used to define classical networks. In the present paper we study how dynamics of the classical networks are reflected in the properties of the QL states.
@@
The physical motivation for the study is to investigate, using numerical studies, how the QL states evolve as a result of evolution of the underlying classical network. We explore the way synchronization and de-synchronization of the classical system translates to dynamics in the state space that can be described as purity-preserving dynamics and decoherence of QL states, respectively. We are motivated to ask how similar the properties of QL states are to quantum states. We find that during decoherence of the QL states, the off-diagonal density matrix elements decay and that this decay can be observed in any basis we choose for measurement. More surprisingly, we show that a no-cloning theorem (that is, a state of a QL bit cannot be copied) applies not only to the QL states, but also to the underlying classical system.
@@
The QL bit graphs can represent physical, classical systems. Most obviously these can be networks of phase oscillators, as studied here. The crucial feature of QL bit graphs and the corresponding classical networks is that they encode two states. In our model, each state can be the emergent state of a subgraph of the QL bit graph, or we can choose the superpositions of emergent states as a basis. QL states, more generally, are represented by product graphs. Here we make no assumptions about the physical manifestation of the phase oscillator system underlying the QL bits, but we could imagine it to represent some kind of dynamical structure of fields or particles that collectively form a microscopic QL object[3, 4, 5]. Or, more abstractly, these underlying variables might be thought of as an intuitive way of representing wavelike properties of a quantum particle because they capture the nature of phase. On the other hand, the phase oscillator system can simply be a macroscopic classical network, for which there are numerous examples[6, 7]. Regardless of these possible interpretations, here we study the specific case of coupled phase oscillators serving as a classical system defined by a QL state graph, and a corresponding state space generated by the adjacency matrix of that graph.
@@
In Fig. 1 the correspondence between the classical system of phase oscillators and the graph and the QL state space is shown. The graph plays a special dual role by directing design of the classical system and defining the state space. The QL state space, specifically, comprises the set of states, projected out of the eigenvectors of the adjacency matrix, that are indicated by the emergent states. These emergent states are highlighted by the box over the spectrum. The interplay between the classical system and these states comes about in the the present work by the evolution of the oscillator phases. We say more about how we configure the classical system in relation to the state space in the next section of the paper.
We develop the idea that there are two viewpoints for our system. First, we can observe the nonlinear dynamics of the complex classical network either at a very detailed classical level, or by considering a global order parameter. Second, we can study the state space in the reduced representation of the emergent QL states. The graph model for QL states[1] connects these two viewpoints.
Two questions need clarification. First, if we represent the QL graph as a network (in the present work, a network of oscillators), what is the difference between the graph and the classical network, because they appear to look the same? The graph encodes rules, according to those rules the classical network is constructed. We can copy the structure of the graph to the network, like we do here. But note that the network needs other features including an assignment of frequencies to the vertices and, possibly, weighting of the edges (couplings). However, we can abstract the design rules and choose a classical system with many more vertices or a continuum of vertices—as long as the structure of each subgraph is $d$-regular then the spectrum will be the same as the graph. Notably, the classical system does not have a state space, but the graph can generate the state space we need via its adjacency matrix.
Second, this state space is defined by the adjacency matrix associated with the graph, so what is the difference between the graph and the QL states? The states are generated by the graph via its adjacency matrix, but they do not uniquely define the graph and its physical properties, as noted below. If we know a gate operation or unitary transformation we want to perform on the QL states, we can find an operation on the QL graph that produces the desired transformation in the state space[8]. However, it is not always obvious how the graph should be transformed. For example, a unitary map on a state does not correspond to an obvious structural change, like an isomophism on the graph. The unitary map operates on the emergent QL states, like those defined in Sec. IV.B, and is defined by a linear map preserving inner product. The graph isomorphism is a bijective map between vertex sets that preserves adjacency and thereby produces two distinct graphs with the same properties. As long as we concomitantly redefine the projections to the QL basis, the states are unchanged.
What seems particularly interesting about the parallel view of the same dynamics, is that it offers a way to think about the underlying classical variables that control properties of the QL state space—those variables can be hidden if we imagine we can only observe the state space and do not know much, or anything, about the classical system that is controlling the states. However, they are not necessarily deterministic, depending how we define the averaging over the system preparation (see the next section).
II How graphs map classical variables to the state space
To start, we briefly provide some background. A graph $G(n,m)$, or simply $G$, comprises $n$ vertices and a set of $m$ edges that connect pairs of vertices. The size of a graph or subgraph, that is, the number of vertices, is written $|G|$. A graph has two important features. First, the structure of the graph—determined by the pattern of edges—plays a key role in the classical dynamical evolution of a system represented by the graph. That is because the structure of couplings between phase oscillators in the network plays an important role in determining phase synchronization[9, 10, 11, 12]. Second, any graph can be characterized by its spectrum and associated eigenstates, but not uniquely[13, 14]. In the present context, for instance, it is possible to find various different (not isomorphic) random graphs, that are on average $d$-regular, with the same spectrum.
Special graphs, called expander graphs, have a spectrum comprising a distinguished emergent state—a state that is clearly separated from other states in the spectrum. The emergent state of an expander graph is the resource we use to construct the QL state space. Two expander graphs are bound by connecting edges to produce a two-state system that we call the QL bit[1]. These QL bit graphs are combined using a graph operation—the Cartesian product of graphs—to generate a new, exponentially larger, graph. The states of this product graph are linear combinations of tensor products of the states of the QL bits[2]. For background see ref [21].
We consider networks of phase oscillators for the classical system, so the implications of phase in the state vectors comes to the forefront. For instance, we show in the results of this work how phase differences lead to decoherence of states when the oscillator network fails to remain synchronized. We thus find that collapse of superposition is possible in an isolated system because of the way the states are generated by a system that cannot lock the phases of the states. We further show how the preferred basis (or pointer states[22]) are defined by the way we observe the QL states.
@@ II How graphs map classical variables to the state space
Any QL state space is mirrored by a corresponding graph in the sense that the graph dictates certain design rules that the classical system must satisfy. For example, we require subsystems that produce emergent states and these subsystems are coupled in a particular way so that correlations are built into the system that match those defining a quantum state space. The classical system should not be confused with an arbitrary classical system. For instance, if the states obey a no-cloning theorem (which they do), then this cannot be circumvented by an allowed transformation of the classical network (see Sec. V.F).
@@
In this report, we study how dynamics of the classical system compares to, and governs, dynamics in the QL state space. The basis for the study are the nonlinear dynamics of a large, complex classical network of phase oscillators. We ask, how are these dynamics reflected in the reduced space of QL states? The work substantiates the claim made in our prior work that QL states can be robust to disorder and resist decoherence, which is demonstrated here by synchronization conditions for the classical network. In contrast, when the classical network does not synchronize, then an initially prepared synchronized system produces QL states that become more mixed with time—they show decoherence. It is shown how we can observe that decoherence in any basis we choose according to our measurement. In the case of QL states undergoing decoherence, we can think of the poorly-synchronized classical network as a kind of intrinsic bath.
III Synchronization
@@
where $10<\mu<<$ 1 is a small parameter that determines how close the system is to a linear conservative system. Now define polar coordinates
@@
$x=\rho\cos(\omega_{0}t+\theta)$ (4)
$y=-\rho\omega_{0}\sin(\omega_{0}t+\theta)$ (5)
@@
$\dot{\rho}\cos(\phi)-\rho\sin(\phi)\dot{\theta}=0$ (6)
$-\dot{\rho}\omega_{0}\sin(\phi)-\rho\omega_{0}\cos(\phi)\dot{\theta}=\mu f(x,y)$ (7)
where $\phi=\omega_{0}t+\theta$, $x=\rho\cos\phi$, and $y=-\rho\omega_{0}\sin\phi$. The solutions are written
$\dot{\rho}=\frac{-\mu f(x,y)\sin\phi}{\omega_{0}}$ (8)
$\dot{\phi}=\omega_{0}-\frac{\mu f(x,y)\cos\phi}{\rho\omega_{0}}$ (9)
Typically $\rho$ is slowly varying compared to $\phi$. The nonlinearity can arise for various reasons, and is reflected in $f(x,y)$, case by case. For example, the pendulum clock is subjected to regular, short pushes that counteract dissipative losses of the pendulum. Electronic oscillator circuits incorporate a nonlinear element that, with the right conditions, enables the circuit to produce sustained oscillations because of feedback between the oscillatory system and the nonlinear element.
@@
where $\nu_{i}$ is the frequency of the oscillator at node $i$, $\theta_{i}$ is the oscillator phase defined in terms of accumulated phase and an offset $\theta_{i}(t)=\epsilon_{i}t+\phi_{i}$. $K$ is the coupling value, $a_{ij}$ (with value 0 or 1) are taken from the adjacency matrix of the graph. The $a_{ij}$ represent the node connectivities in the network. The coupling is renormalized by the factor $1/N$ so that networks are size-intensive with respect to the lowest eigenvalue. Notice that the nonlinearity comes about because of the terms $f(\theta_{j}-\theta_{i})$, where $f(\dots)=\sin(\dots)$.
@@
order $\text{order parameter}=\text{Re}\Big{(}\frac{1}{N}\sum_{i=1}^{N}e^{i\theta_{i}}\Big{)}.$ (11)
@@
IV Phase oscillator synchronization and dynamics in the state space
IV.1 Phase and the adjacency matrix
Here we examine the relationship between, on the one hand, dynamics of a nonlinear classical system, and on the other hand, the evolution of the state space interpreted by the graph representation. The model we use here is to assign each vertex of the graph as a phase oscillator, and the edges indicate couplings between pairs of phase oscillators. Thus we can use the Kuramoto model to explore the synchronization dynamics of the classical network of phase oscillators. The oscillators (vertices) are connected according to the graph model for QL bits and their products. This provides a physical realization of the classical variables associated with the QL states. We concomitantly calculate the states and their dynamics as the network synchronization evolves. It will turn out that this gives insight into, and a tangible physical model for, the implications of phase relationships between the variables underlying QL states.
The eigenvectors we have described in prior work are calculated for the normal adjacency matrices of the graphs (i.e. edges have values ±$\pm 1$). They correspond to the states of perfectly globally-synchronized systems. More generally, however, each oscillator labeled $j$ is associated with a phase $\theta_{j}$. The adjacency matrix $A$ is transformed by these phases according to:

### v4

@@ Abstract
In recent work we have developed a theory for classical systems that mimic the state space of quantum systems. We call the states ‘quantum-like’ (QL) because they show many properties similar to those of a quantum system. Here we study the QL states of a large classical system comprising a network of coupled phase oscillators. We investigate how mixedness of the QL states diminish or increase as the underlying classical system synchronizes or de-synchronizes respectively. This allows study of the interplay between the nonlinear dynamics of the variables of the classical system and the properties of the corresponding QL states. In the limit of a very strongly phase-locked classical network—that is, where couplings between phase oscillators are very large—the state space evolves according to unitary dynamics. When the system is set up so that the classical phase oscillators synchronize with time, we find that the purity measure of the QL states correspondingly increases. Whereas, in cases where the coupling between oscillators is too small to allow the network to synchronize, the classical variables act like a kind of environment that promotes dephasing of initially-prepared superpositions.
@@ I Introduction
In recent work we have described how a classical system can be mapped, via a graph representation, to a state space that has many properties similar to a quantum state space[1, 2]. For that reason, we call it a quantum-like (QL) state space, and refer to the states as QL states. The graphs needed as a basis for these constructions are called QL bit graphs (or simply QL bits). By applying a product operation to QL bit graphs we can produce graphs whose states comprise arbitrary (controllable) superpositions of states in a tensor product basis. These same graphs can also be used to define classical networks. In the present paper we study how dynamics of the classical networks are reflected in the properties of the QL states.
@@
The physical motivation for the study is to investigate, using numerical studies, how the QL states evolve as a result of evolution of the underlying classical network. We explore the way synchronization and de-synchronization of the classical system translates to dynamics in the state space that can be described as purity-preserving dynamics and dephasing of QL states, respectively. Synchronization occurs when the coupling between phase oscillators is sufficiently large that the phases become locked in step; a phenomenon that is extensively studied. When the network is synchronized, we have concomitant emergent QL states. Dephasing of the QL states results from the opposite limit, where couplings are small and the phase oscillators evolve according to their natural frequencies and the network loses phase coherence. As a result, emergent QL states are absent.
@@
The QL bit graphs can represent physical, classical systems. Most obviously these can be networks of phase oscillators, as studied here. The crucial feature of QL bit graphs and the corresponding classical networks is that they encode two states. In our model, each state can be the emergent state of a subgraph of the QL bit graph, or we can choose the superpositions of emergent states as a basis. QL states, more generally, are represented by products of these QL bit graphs. Here we make no assumptions about the physical manifestation of the phase oscillator system underlying the QL bits, but we could imagine it to represent some kind of dynamical structure of fields or particles that collectively form a microscopic QL object[3, 4, 5]. On the other hand, the phase oscillator system can simply be a macroscopic classical network, for which there are numerous examples[6, 7]. Regardless of these possible interpretations, here we study the specific case of coupled phase oscillators serving as a classical system defined by a QL state graph, and a corresponding state space generated by the adjacency matrix of that graph.
II Graphs that generate quantum-like states
Here we give a brief introduction to graphs and to QL states. To see a summary, refer to Fig. 1. Before beginning the explanation of QL graphs and states, it is important to understand how to think about them compared to quantum systems. When we combine QL graphs using the product operation to generate QL states, the QL graphs do not represent quantum systems like electrons or photons, and so on. Rather, they are a physical manifestation of the states of composite quantum systems such as many-electron atoms, entangled photons. This is not something we normally think about. However, we are familiar with the state space represented by these graphs; for instance, states spanned by the Bell states. The Bell states can be generated by diagonalizing a suitable matrix. That matrix can also generate a graph. This is the way of thinking about what the QL graphs mean. A second point to note is that the states considered in the present work are separable, by construction.
Connecting a classical system to a space of states that we normally associate with composite quantum systems is not as heretical as it might, at first, sound. The states of electron spin in an atom or molecule are, via the Dirac identity, given by irreducible representations of the symmetric group. There are obviously many classical systems that, similarly, are governed by the structure of the symmetric group. Schur polynomials, symmetric functions, lattices, and posets provide some examples.
Now we briefly provide some background. A graph $G(n,m)$, or simply $G$, comprises $n$ vertices and a set of $m$ edges that connect pairs of vertices. The size of a graph or subgraph, that is, the number of vertices, is written $|G|$. A graph has two important features. First, the structure of the graph—determined by the pattern of edges—plays a key role in the classical dynamical evolution of a system represented by the graph. That is because the structure of couplings between phase oscillators in the network plays an important role in determining phase synchronization[8, 9, 10, 11]. Second, any graph can be characterized by its spectrum and associated eigenstates, but not uniquely[12, 13]. In the present context, for instance, it is possible to find various different (not isomorphic) random graphs, that are on average $d$-regular, with the same spectrum.
The starting point for graphs that produce QL states is a graph that is a good source of states. By ‘good’, what we mean is that we want a really large and complex graph—imagine part of a neural network in a brain—to have a spectrum that shows one special state that is isolated from all the other many (possibly infinite even) states. That is, we need an emergent state. We therefore focus on special graphs, called expander graphs, that have a spectrum comprising a distinguished emergent state—a state that is clearly separated from other states in the spectrum. The emergent state of an expander graph is the resource we use to construct the QL state space. Two expander graphs are bound by connecting edges to produce a two-state system that we call the QL bit[1]. These QL bit graphs, in turn, are combined using a graph operation—the Cartesian product of graphs—to generate a new, exponentially larger, graph. The states of this product graph are linear combinations of tensor products of the states of the QL bits[2]. For background see ref [20].
Figure 1: A quick-start guide to QL graphs and their states.
An example of a graph that will not work as a reliable source of states is one that represents a system of vertices connected only by nearest-neighbor edges, because as the system gets larger, its spectrum gets more congested. In contrast, we need a graph that guarantees one state is separated from the rest. Underlying this is a topology on the vector space (where our states live) called the Fell topology. If we can construct a graph the exploits that topology, then we find one emergent state that gets ‘pushed away’ from all the other states. An example of such families of graphs are the expander graphs. In particular, optimal expander graphs (i.e. those whose spectra have the best possible gap between the first and second eigenvalues) are known as Ramanujan graphs[21, 22].
It is not strictly necessary that we employ optimal expanders because the relevant, practical graphs will often be small and finite. What we do need is the spectral gap between the emergent state and all other states. It is also desirable to have the robustness of an expander graph, that is, that it doesn’t matter too much how we throw the graph together, if we follow a couple of simple rules it will work out. In fact, we could as well use an Erdös-Renyi random graph[19, 18] with sufficient edges and not worry at all about how we wire up the vertices, it will work[23]. Nevertheless, it is nice to use expander families because we might need to use their asymptotic properties for certain proofs or to take a continuum limit.
We consider networks of phase oscillators for the classical system, so the implications of phase in the state vectors comes to the forefront. For instance, we show in the results of this work how phase differences lead to dephasing of states when the oscillator network fails to remain synchronized. We thus find that collapse of a QL superposition is possible in an isolated network because when the oscillator phases are not locked together, the emergent state is lost.
III Purpose of the paper
@@
In Fig. 2 the correspondence between a classical system of phase oscillators and the graph and the QL state space is shown. The classical system should not be confused with an arbitrary classical system. The graph plays a special dual role by directing design of a special classical system and, in the other direction, the graph defines a QL state space that is paired to the special classical system. What this means is that the graph dictates certain design rules that the classical system must satisfy. Obviously this is crucial because there are infinite classical systems that cannot be mapped to a QL state space. We need to pinpoint those rare kinds of classical systems that can generate QL states. For example, we require subsystems that produce emergent states and these subsystems are coupled in a particular way so that correlations are built into the system that match those defining a quantum state space.
The QL state space, specifically, comprises the set of states, projected out of the eigenvectors of the adjacency matrix, that are indicated by the emergent states. These emergent states are highlighted by the box over the spectrum. The interplay between the classical system and these states comes about in the present work by the evolution of the oscillator phases. We use this model to develop the idea that there are two viewpoints for our system. First, we can observe the nonlinear dynamics of the complex classical network either at a very detailed classical level, or by considering a global order parameter. Second, we can study the state space in the reduced representation of the emergent QL states. The graph model for QL states[1] connects these two viewpoints.
Two questions need clarification. First, if we represent the QL graph as a network (in the present work, a network of oscillators), what is the difference between the graph and the classical network, because they appear to look the same? The graph encodes rules, according to those rules the classical network is constructed (as explained above). We can copy the structure of the graph to the network, like we do here. But note that the network needs other features including an assignment of frequencies to the vertices and, possibly, weighting of the edges (couplings). However, we can abstract the design rules and choose a classical system with many more vertices or a continuum of vertices—as long as the structure of each subgraph is $d$-regular then the spectrum will be the same as the graph. Notably, the classical system does not have a state space, but the graph can generate the state space we need via its adjacency matrix.
Second, this state space is defined by the adjacency matrix associated with the graph, so what is the difference between the graph and the QL states? The states are generated by the graph via its adjacency matrix, but they do not uniquely define the graph and its physical properties, as noted below. If we know a gate operation or unitary transformation we want to perform on the QL states, we can find an operation on the QL graph that produces the desired transformation in the state space[24]. However, it is not always obvious how the graph should be transformed. For example, a unitary map on a state does not correspond to an obvious structural change, like an isomophism on the graph. The unitary map operates on the emergent QL states and is defined by a linear map preserving inner product. The graph isomorphism is a bijective map between vertex sets that preserves adjacency and thereby produces two distinct graphs with the same properties. As long as we concomitantly redefine the projections to the QL basis, the states are unchanged.
What seems particularly interesting about the parallel view of the same dynamics, is that it offers a way to think about the underlying classical variables that control properties of the QL state space—those variables can be hidden if we imagine we can only observe the state space and do not know much, or anything, about the classical system that is controlling the states. However, they are not necessarily deterministic, depending how we define the averaging over the system preparation.
@@
In this report, we study specifically how dynamics of the classical system compares to, and governs, dynamics in the QL state space. The basis for the study are the nonlinear dynamics of a large, complex classical network of phase oscillators. We ask, how are these dynamics reflected in the reduced space of QL states? The work substantiates the claim made in our prior work that QL states can be robust to disorder and resist decoherence. That is shown when the couplings between phase oscillators are sufficiently strong so that the classical network synchronizes despite an initial random distribution of phases. In contrast, when the classical network does not synchronize, then an initially prepared system produces QL states that become more mixed with time—they show dephasing. In the case of QL states undergoing dephasing, we can think of the poorly-synchronized classical network as a kind of intrinsic bath that mixes the pure emergent states.
IV Synchronization
@@
where $0<\mu<<1$ is a small parameter that determines how close the system is to a linear conservative system. Now define polar coordinates
@@
$x=\rho\cos(\omega_{0}t+\phi)$ (4)
$y=-\rho\omega_{0}\sin(\omega_{0}t+\phi)$ (5)
@@
$\dot{\rho}\cos(\theta)-\rho\sin(\theta)\dot{\phi}=0$ (6)
$-\dot{\rho}\omega_{0}\sin(\theta)-\rho\omega_{0}\cos(\theta)\dot{\phi}=\mu f(x,y)$ (7)
where $\theta=\omega_{0}t+\phi$, $x=\rho\cos\theta$, and $y=-\rho\omega_{0}\sin\theta$. The solutions are written
$\dot{\rho}=\frac{-\mu f(x,y)\sin\theta}{\omega_{0}}$ (8)
$\dot{\theta}=\omega_{0}-\frac{\mu f(x,y)\cos\theta}{\rho\omega_{0}}$ (9)
Typically $\rho$ is slowly varying compared to $\theta$. The nonlinearity can arise for various reasons, and is reflected in $f(x,y)$, case by case. For example, the pendulum clock is subjected to regular, short pushes that counteract dissipative losses of the pendulum. Electronic oscillator circuits incorporate a nonlinear element that, with the right conditions, enables the circuit to produce sustained oscillations because of feedback between the oscillatory system and the nonlinear element.
@@
[…]
@@
$\text{order parameter}=\text{Re}\Big{(}\frac{1}{N}\sum_{i=1}^{N}e^{i\theta_{i}}\Big{)}.$ (11)
@@
V Phase oscillator synchronization and dynamics in the state space
V.1 Phase and the adjacency matrix
Here we examine the relationship between, on the one hand, dynamics of a nonlinear classical system, and on the other hand, the evolution of the state space interpreted by the graph representation. The model we use is to assign each vertex of the graph as a phase oscillator, and the edges indicate couplings between pairs of phase oscillators. Thus we can use the Kuramoto model to explore the synchronization dynamics of the classical network of phase oscillators. The oscillators (vertices) are connected according to the graph model for QL bits and their products. This provides a physical realization of the classical variables associated with the QL states. We concomitantly calculate the states and their dynamics as the network synchronization evolves. It will turn out that this gives insight into, and a tangible physical model for, the implications of phase relationships between the variables underlying QL states.
Specifically, what we do in the rest of the paper is to take a very large network of oscillators, organized according to a graph for a composite QL state system formed by the graph product of two QL bits. We propagate the nonlinear dynamics of this network according to the Kuramoto model. Now every vertex $v_{i}$ in this large graph is associated with the phase $\theta_{i}$ of the corresponding phase oscillator in the network. What we want to study is how the corresponding QL state—the one state with highest eigenvalue in the way we write the adjacency matrix—evolves. Notice that we write the Kuramoto equation so that the coupling term is preceded by a minus sign, so this state that has the largest eigenvalue for the adjacency matrix is the ground state of the classical model. To study the QL state under the Kuramoto dynamics, we need to modify the adjacency matrix so that it includes the evolving phases that are associated with the graph vertices. We now show that we do that by using a unitary transformation at each time step.
This result is important because it gives the reason that the nonlinear dynamics of the Kuramoto model are represented by linear dynamics (but not necessarily unitary dynamics) in the state space. The point is confusing; how are the nonlinear dynamics of the phase oscillator model represented by linear evolution in the QL state space? The phases in Kuramoto model evolve nonlinearly, but they transform the adjacency matrix through a unitary operator. By definition this preserves inner products, so the evolution of the state happens by a sequence of linear maps. We emphasize that statement below in more detail. The result is particular for the phase oscillator model and might not be generalizable. However, it is known that the nonlinear action of groups can be linearized by using a matrix representation of the group, which necessarily transforms according to linear algebra. This is more-or-less what we are trying to accomplish through the graph representation of the classical system.
The eigenvectors we have described in prior work, the QL states, are calculated for the normal adjacency matrices of the graphs (i.e. edges have values ±1$\pm$ 1± 1). They correspond to the states of perfectly globally-synchronized systems. More generally, however, each oscillator labeled $j$ is associated with a phase $\theta_{j}$. The adjacency matrix $A$ is transformed by these phases according to: